# A unified approach to hypergeometric class functions

### v1

@@ 4.3 Power symmetry
$z^{m}\Big{(}{\mathcal{C}}(\sigma,\kappa)+\omega\Big{)}z^{-m}={\mathcal{C}}(\sigma,\kappa^{\circ})+\omega^{\circ}.$ (4.12)
@@
$z^{m}\Big{(}{\mathcal{C}}(\sigma,\kappa)+\omega\Big{)}z^{-m}$ (4.14)
@@
$=$ ${\mathcal{C}}(\sigma,\kappa^{\circ})+\omega^{\circ}.$ (4.17)
Note that $\kappa^{\circ\circ}=\kappa$ and $\omega^{\circ\circ}=\omega$. Thus we obtain a ${\mathbb{Z}}_{2}$ symmetry of hypergeometric class equations, different from the basic symmetry of Thm 3.10.
As a consequence of (4.12), both $\mathbf{F}(\sigma,\kappa,\omega;z)$ and $z^{-m}{\mathbf{F}}(\sigma,\kappa^{\circ},\omega^{\circ};z)$ are annihilated by ${\mathcal{C}}(\sigma,\kappa)+\omega$. If $m\not\in{\mathbb{Z}}$ they form a basis of solutions of the corresponding equation. The situation for $m\in{\mathbb{Z}}$ will be discussed in Section 8 about the degenerate case.
@@ 4.4 Inversion symmetry
$-z^{-\zeta+1}\Big{(}{\mathcal{C}}(\sigma,\kappa;z,\partial_{z})+\omega\Big{)}z^{\zeta}={\mathcal{C}}(\sigma^{\triangle},\kappa^{\triangle};w,\partial_{w})+\omega^{\triangle}$ (4.23)
@@
${\mathcal{C}}(\sigma,\kappa)+\omega\quad\text{ annihilates }\quad z^{-\zeta}\mathbf{F}(\sigma^{\triangle},\kappa^{\triangle},\omega^{\triangle};-z^{-1}).$ (4.24)
@@
$-z\Big{(}{\mathcal{C}}(\sigma,\kappa;z,\partial_{z})+\omega\Big{)}$ (4.25)
@@
Note that $\sigma^{\triangle\triangle}=\sigma$, $\kappa^{\triangle\triangle}=\kappa$, $\omega^{\triangle\triangle}=\omega$. Hence the inversion symmetry generates a group isomorphic to ${\mathbb{Z}}_{2}$.
@@
It is easy to see that the substitution $z\mapsto\frac{1}{w}$ also generates a symmetry of hypergeometric class equations. This follows immediately from Theorem 4.24 and the trivial fact that $z\mapsto-z$ leads to a symmetry as well.
@@ 5 Integral representations
Hypergeometric class functions possess useful integral representations. Most of them have the form of an Euler transform of an elementary function. In Subsection 5.1 we will show how to derive these representaions in a unified way. Note that typically for a given equation one can choose various contours of integration, obtaining various solutions.
@@ 5.1 Euler transforms
${\cal C}(\sigma,\kappa_{n})+\omega_{n}$ (5.6)
@@
$\partial_{z}f_{n}(z)$ $=(n+1)f_{n+1}(z)$ (5.8)
$\left(\sigma(z)\partial_{z}+\kappa_{n+1}(z)\right)f_{n+1}(z)$ $=-\left(\kappa_{0}^{\prime}+\frac{\sigma^{\prime\prime}}{2}n\right)f_{n}(z)$ (5.9)
@@
$\begin{array}[]{rl}\partial_{z}\mathbf{f}_{n}(z)&=\mathbf{f}_{n+1}(z)\\[5.69054pt] \left(\sigma(z)\partial_{z}+\kappa_{n+1}(z)\right)\mathbf{f}_{n+1}(z)&=-(n+1)\left(\kappa_{0}^{\prime}+\frac{\sigma^{\prime\prime}}{2}n\right)\mathbf{f}_{n}(z)\end{array}$ (5.19)
@@ 5.2 Laplace integrals
${\cal C}(\sigma,\kappa_{n})+\omega_{n}$ (5.24)
@@ 6 Applications case by case
In this section we consider all five nontrivial types of hypergeometric equations, listed in the introduction. We describe those of their properties that are direct consequences of the unified theory, discussed in the previous sections. We try to follow the same pattern. Here is the list of items that we will give for each hypergeometric type, if available:
  • Operator that generates the equation, that is ${\mathcal{C}}(\sigma,\kappa)+\omega$, expressed in the traditional form.
@@
  • Standard solution annihillated by ${\mathcal{C}}(\sigma,\kappa)+\omega$ (typically a special case of the unified hypergeometric function (4.4), but for the Hermite equation we need to make an exception).
@@
We will often use the Pochhammer symbols
@@ 6.2 The ${}_{1}F_{1}$ equation
$z^{a+1}{\cal F}(a;c;z,\partial_{z})z^{-a}={\cal F}(a,1+a-c;_{;}w,\partial_{w}),\qquad w=-z^{-1}.$
@@
$\frac{1}{2\pi{\rm i}}\int\limits_{]-\infty,(0,z)^{+},-\infty[}t^{a-c}{\rm e}^{t}(t-z)^{-a}{\rm d}t$ $=$ ${\bf F}(a;c;z),$
@@
$\int\limits_{[1,+\infty[}{\rm e}^{\frac{z}{t}}t^{-c}(t-1)^{c-a-1}{\rm d}t$ $=$ $\Gamma(a)\Gamma(c-a)\mathbf{F}(a;c;z),$
@@ 6.3 The ${}_{2}F_{0}$ equation
where $|\arg c-\pi|<\pi-\epsilon$, $\epsilon>0$. It extends to an analytic function on the universal cover of \$\0{\mathbb{C}}\backslash\{0\}$ with a branch point of an infinite order at 0. It is annihilated by ${\cal F}(a,b;-)$.
@@
$F(a,b;-;z)\sim\sum_{j=0}^{\infty}\frac{(a)_{j}(b)_{j}}{j!}z^{j},\ |\arg z-\pi|<\pi-\epsilon.$
@@ 6.4 The ${}_{0}F_{1}$ equation
Basic and power symmetry:
@@
$F(c;z):=\sum_{j=0}^{\infty}\frac{1}{(c)_{j}}\frac{z^{j}}{j!}$
@@
$\partial_{z}\mathbf{F}(c;z)$ $=\mathbf{F}(c+1;z)$ (6.25)
@@
$\frac{1}{2\pi{\rm i}}\int\limits_{]-\infty,0^{+},-\infty[}{\rm e}^{t}{\rm e}^{\frac{z}{t}}t^{-c}{\rm d}t$ $=$ ${\bf F}(c;z),\ \ \ \ \mathrm{Re}z>0;$
@@ 6.5 Hermite equation
Integral representation for $z\not\in]-\infty,0]$. For $-\frac{1}{2}<{\mathrm{Re}}\omega$:
$\int\limits_{0}^{\infty}{\rm e}^{-t^{2}-2tz}t^{a-1}{\rm d}t$ $=$ $2^{-a}\Gamma(a)S(a;z),$
@@ 6.6 Lie algebras of symmetries type by type
Miller’s Lie algebra is responsible only for one pair of recurrence relations of hypergeometric class functions. It leads also to a group of discrete symmetries ${\mathbb{Z}}_{2}$. Some of types from the hypergeometric class possess larger sets of recurrence relations, and larger groups of discrete symmetries, which were described e.g. in [De1, De2]. They also possess larger Lie algebras of symmetries, described e.g. in [De2].
In the following table we list all nontrivial types of hypergeometric class equations. We include the Gegenbauer equation, which is the subtype of the ${}_{2}F_{1}$ type with the mirror symmetry $z\mapsto-z$. In this paper there has been no need to consider it separately, however it has special symmetry properties.
The second and third column are based on Section 2 of this paper. The fourth and fifth column are based on [De2]. They describe a more complete Lie algebra of symmetries and the corresponding group of discrete symmetries.
@@
Above, $sch(n,{\mathbb{C}})$ denotes the complex Schrödinger Lie algebra, the Lie algebra of generalized symmetries of the heat equation in dimension $n$—see [De2].
The Lie algebra of the fourth column always contains the corresponding Miller’s Lie algebra (possibly, without the trivial term $\oplus{\mathbb{C}}$). Similarly, the group from the fifth column always contains the Weyl group of Miller’s Lie algebra.
@@ 7 Hypergeometric polynomials
The following well-known and easy proposition shows that hypergeometric class operators have many polynomial eigenfunctions. In some sense, it is the property that characterizes this class among all second order differential operators.
@@
Let $\sigma(z)$, $\tau(z)$, $\mu(z)$ be arbitrary functions. Let $P_{n}(z)$, $n=0,1,\dots$, be polynomials such that $\deg P_{n}=n$ and $\eta_{n}\in{\mathbb{C}}$. Suppose that
$\left(\sigma(z)\partial_{z}^{2}+\tau(z)\partial_{z}+\mu(z)+\eta_{n}\right)P_{n}(z)=0,\ \ n=0,1,\dots$
Then $\sigma(z)$ is a polynomial of degree at most $2$, $\tau(z)$ is a polynomial of degree at most 1 and $\mu(z)$ a polynomial of degree 00 (a complex number). Thus, replacing $\mu+\eta_{n}$ with $\eta_{n}$, without limiting the generality we can assume that $\mu=0$. We then have
$\begin{array}[]{l}n(n-1)\frac{\sigma^{\prime\prime}}{2}+n\tau^{\prime}+\eta_{n}=0.\end{array}$ (7.1)
Here is another simple fact about polynomial solutions of hypergeometric class equations:
@@
Let $\deg\sigma\leq 2$ and $\deg\tau\leq 1$ and $\eta\in{\mathbb{C}}$ and
$k\frac{\sigma^{\prime\prime}}{2}+\tau^{\prime}\neq 0,\quad k=0,1,2,\dots.$ (7.2)
Then the space of polynomial solutions of
$\left(\sigma(z)\partial_{z}^{2}+\tau(z)\partial_{z}+\eta\right)P(z)=0$ (7.3)
is at most $1$-dimensional.
Proof.  Suppose the space of polynomial solutions of (7.3) is 2-dimensional. We can assume that the degrees of these solutions are $0\leq n_{1}<n_{2}$. By (7.1),
@@
The following theorem shows how to construct hypergeometric class polynomials and describes some of their properties, which are easy to describe in a unified fashion. As usual, $\sigma$ denotes a polynomial of degree $\leq 2$, $\kappa$ a polynomial of degree $\leq 1$ and
1. and 2. of the following theorem is just a reformulation of Propositions 7.1 and 7.2, where instead of parameters $\tau,\eta$ we use $\kappa,\omega$.
@@
    00 0 $=\Big{(}{\cal C}(\sigma,\kappa)+\omega\Big{)}v(z)$ (7.6)
    $=\Big{(}\sigma(z)\partial_{z}^{2}+(\sigma^{\prime}(z)+\kappa(z))\partial_{z}+\frac{\kappa^{\prime}}{2}+\omega\Big{)}v(z).$ (7.7)
@@
    00 0 $=\Big{(}{\cal C}(\sigma,\kappa)-n(n+1)\frac{\sigma^{\prime\prime}}{2}-(n+\frac{1}{2})\kappa^{\prime}\Big{)}P_{n}(\sigma,\kappa;z)$ (7.12)
@@
Proof.  The integral representation (7.11) satisfies the assumptions of Theorem 5.1 about Euler transforms. This implies (7.12) holds. (7.10), usually called the Rodriguez formula, follows immediately from (7.11). $\Box$
Consider the descending ladder starting at $\kappa_{0}(z):=\kappa(z)$, $\omega_{0}=-\frac{\kappa^{\prime}}{2}$, that is
@@
$\Big{(}{\cal C}(\sigma,\kappa_{-n})+\omega_{-n}\Big{)}P_{n}(\sigma,\kappa_{-n};z)$ $=0,$ (7.15)

### v2

@@ 4.3 Power symmetry
$z^{m}\Big{(}{{\cal H}}(\sigma,\kappa)+\omega\Big{)}z^{-m}={{\cal H}}(\sigma,\kappa^{\circ})+\omega^{\circ}.$ (4.12)
@@
$z^{m}\Big{(}{{\cal H}}(\sigma,\kappa)+\omega\Big{)}z^{-m}$ (4.14)
@@
$=$ ${{\cal H}}(\sigma,\kappa^{\circ})+\omega^{\circ}.$ (4.17)
Note that $\kappa^{\circ\circ}=\kappa$ and $\omega^{\circ\circ}=\omega$. Thus we obtain a ${\mathbb{Z}}_{2}$ symmetry of the hypergeometric class. It is different from the basic symmetry of Thm 3.10.
As a consequence of (4.12), both $\mathbf{F}(\sigma,\kappa,\omega;z)$ and $z^{-m}{\mathbf{F}}(\sigma,\kappa^{\circ},\omega^{\circ};z)$ are annihilated by ${{\cal H}}(\sigma,\kappa)+\omega$. If $m\not\in{\mathbb{Z}}$ they form a basis of solutions of the corresponding equation. The situation for $m\in{\mathbb{Z}}$ will be discussed in Section 8 about the degenerate case.
@@ 4.4 Inversion symmetry
$-z^{-\zeta+1}\Big{(}{{\cal H}}(\sigma,\kappa;z,\partial_{z})+\omega\Big{)}z^{\zeta}={{\cal H}}(\sigma^{\triangle},\kappa^{\triangle};w,\partial_{w})+\omega^{\triangle}$ (4.23)
@@
${{\cal H}}(\sigma,\kappa)+\omega\quad\text{ annihilates }\quad z^{-\zeta}\mathbf{F}(\sigma^{\triangle},\kappa^{\triangle},\omega^{\triangle};-z^{-1}).$ (4.24)
@@
$-z\Big{(}{{\cal H}}(\sigma,\kappa;z,\partial_{z})+\omega\Big{)}$ (4.25)
@@
Note that $\sigma^{\triangle\triangle}=\sigma$, $\kappa^{\triangle\triangle}=\kappa$, $\omega^{\triangle\triangle}=\omega$. Hence the inversion symmetry generates a group isomorphic to ${\mathbb{Z}}_{2}$ acting on the hypergeometric class.
@@
It is easy to see that the substitution $z\mapsto\frac{1}{w}$ also generates a symmetry of hypergeometric class equations. This follows immediately from Theorem 4.24 and the trivial fact that $z\mapsto-z$ leads to a symmetry of the hypergeometric class as well.
@@ 5 Integral representations
Hypergeometric class functions possess useful integral representations. Most of them have the form of an Euler transform of an elementary function. In Subsection 5.1 we will show how to derive these representations in a unified way. Note that typically for a given equation one can choose various contours of integration, obtaining various solutions.
@@ 5.1 Euler transforms
${{\cal H}}(\sigma,\kappa_{n})+\omega_{n}$ (5.6)
@@
$\partial_{z}f_{n}(z)$ $=(n+1)f_{n+1}(z),$ (5.8)
$\left(\sigma(z)\partial_{z}+\kappa_{n+1}(z)\right)f_{n+1}(z)$ $=-\left(\kappa_{0}^{\prime}+\frac{\sigma^{\prime\prime}}{2}n\right)f_{n}(z).$ (5.9)
@@
$\begin{array}[]{rl}\partial_{z}\mathbf{f}_{n}(z)&=\mathbf{f}_{n+1}(z),\\[5.69054pt] \left(\sigma(z)\partial_{z}+\kappa_{n+1}(z)\right)\mathbf{f}_{n+1}(z)&=-(n+1)\left(\kappa_{0}^{\prime}+\frac{\sigma^{\prime\prime}}{2}n\right)\mathbf{f}_{n}(z).\end{array}$ (5.19)
@@ 5.2 Laplace integrals
${{\cal H}}(\sigma,\kappa_{n})+\omega_{n}$ (5.24)
@@ 6 Applications case by case
Hypergeometric class operators can be divided into several types. By an affine transformation $z\mapsto az+b$ and division by a constant, an operator of each type can be reduced to its normal form. There are 5 nontrivial types, with normal forms listed in the introduction. For instance, if $\sigma$ has two distinct roots, then the equation belongs to the ${}_{2}F_{1}$ type, if it has a double root, it belongs to the ${}_{2}F_{0}$ type, etc. The hypergometric class contains also 4 “trivial types”, which can be solved in an elementary way. All these 9 types are listed in the table in Appendix B.
Strictly speaking, this table is devoted to types of the Riemann class, which is larger than the hypergeometric class. Therefore, this table contains 10 types. It includes one additional type: the Airy operator, which cannot be reduced to the hypergeometric class. However, all the 9 other types of the Riemann class are represented in the hypergeometric class.
In this section we describe the properties of all five nontrivial types of the hypergeometric class that are direct consequences of the unified theory, discussed in the previous sections. We try to follow the same pattern. Here is the list of items that we will give for each hypergeometric type, if available:
  • Operator that generates the equation, that is ${{\cal H}}(\sigma,\kappa)+\omega$, expressed in the traditional form.
@@
  • Standard solution annihillated by ${{\cal H}}(\sigma,\kappa)+\omega$ (typically a special case of the unified hypergeometric function (4.4), but for the Hermite equation we need to make an exception).
@@
We will often use the Pochhammer symbol
@@ 6.2 The ${}_{1}F_{1}$ equation
$z^{a+1}{\cal F}(a;c;z,\partial_{z})z^{-a}={\cal F}(a,1+a-c;{-};w,\partial_{w}),\qquad w=-z^{-1}.$
@@
$\frac{1}{2\pi{\rm i}}\int\limits_{]-\infty,(0,z)^{+},-\infty[}t^{a-c}{\rm e}^{t}(t-z)^{-a}{\rm d}t$ $=$ ${\bf F}(a;c;z);$
@@
$\int\limits_{[1,+\infty[}{\rm e}^{\frac{z}{t}}t^{-c}(t-1)^{c-a-1}{\rm d}t$ $=$ $\Gamma(a)\Gamma(c-a)\mathbf{F}(a;c;z).$
@@ 6.3 The ${}_{2}F_{0}$ equation
where $|\arg c|>\epsilon$, $\epsilon>0$. It extends to an analytic function on the universal cover of \$\0{\mathbb{C}}\backslash\{0\}$ with a branch point of an infinite order at 0. It is annihilated by ${\cal F}(a,b;-)$.
@@
$F(a,b;-;z)\sim\sum_{j=0}^{\infty}\frac{(a)_{j}(b)_{j}}{j!}z^{j},\ |\arg z|>\epsilon.$
@@
The function ${}_{2}F_{0}$ is in our opinion insufficiently known. Therefore, we devote Appendix A to a derivation of some of its properties.
Note that the equivalence of ${}_{2}F_{0}$ and ${}_{1}F_{1}$ equations is helpful in deriving properties of Subsection 6.2 from Subsection 6.3. For instance, the integral representation (6.18) can be deduced from the integral representation (6.12) and the inversion symmetry. However, the relationship between some of the properties is not so straightforward. For instance, the basic recurrence relations for ${}_{2}F_{0}$ equations do not follow directly from the basic recurrence relations for the ${}_{1}F_{1}$ equations. Besides ${}_{2}F_{0}$ equation does not possess the power symmetry, even though the ${}_{1}F_{1}$ equation has it. Instead, the ${}_{2}F_{0}$ equation is symmetric with respect to the parameter interchange $a\leftrightarrow b$.
@@ 6.4 The ${}_{0}F_{1}$ equation
Basic symmetry $=$ power symmetry:
@@
$F(c;z):=\sum_{j=0}^{\infty}\frac{1}{(c)_{j}}\frac{z^{j}}{j!}.$
@@
$\partial_{z}\mathbf{F}(c;z)$ $=\mathbf{F}(c+1;z),$ (6.25)
@@
$\frac{1}{2\pi{\rm i}}\int\limits_{]-\infty,0^{+},-\infty[}{\rm e}^{t}{\rm e}^{\frac{z}{t}}t^{-c}{\rm d}t$ $=$ ${\bf F}(c;z),\ \ \ \ \mathrm{Re}z>0.$
@@ 6.5 Hermite equation
Integral representation for $z\not\in]-\infty,0]$. For $0<{\mathrm{Re}}a$:
$\int\limits_{0}^{\infty}{\rm e}^{-t^{2}-2tz}t^{a-1}{\rm d}t$ $=$ $2^{-a}\Gamma(a)S(a;z);$
@@ 6.6 Lie algebras of symmetries type by type
Miller’s Lie algebra is responsible only for one pair of recurrence relations of hypergeometric class functions. Some of types from the hypergeometric class possess larger sets of recurrence relations, which were described e.g. in [De1, De2]. These recurrence relations can be interpreted in terms of roots of larger Lie algebras of symmetries, described e.g. in [De2].
Miller’s Lie algebra possesses a Weyl group isomorphic to ${\mathbb{Z}}_{2}$. Some types of hypergeometric class equations possess larger groups of discrete symmetries.
In the following table we list all nontrivial types of hypergeometric class equations. We include the Gegenbauer equation
$\Big{(}(1-w^{2})\partial_{w}^{2}-2(1+\alpha)w\partial_{w}+\lambda^{2}-\big{(}\alpha+\tfrac{1}{2}\big{)}^{2}\Big{)}f(w)=0,$ (6.35)
treating it as a seperate type. Strictly speaking, the Gegenbauer type is contained in the ${}_{2}F_{1}$ type. In fact, the transformation $w\mapsto\frac{1-w}{2}$ transforms the Gegenbauer equation into a special case of the ${}_{2}F_{1}$ equation. The Gegenbauer equation has the special property of the mirror symmetry $w\mapsto-w$. In the remaining part of this paper there is no need to consider it separately, however it has special symmetry properties.
The second and third column are based on Section 2 of this paper. In particular, the second column can be compared with (2.18)–(2.20).
The fourth and fifth column are based on [De2]. They describe a more complete Lie algebra of symmetries and the corresponding group of discrete symmetries.
@@
Above, $sch(n,{\mathbb{C}})$ denotes the complex Schrödinger Lie algebra, the Lie algebra of generalized symmetries of the heat equation in dimension $n$—see e.g. [De2].
The group from the fifth column always contains various symmetries described in our paper. First of all, all types possess the basic symmetry, which is the generator of the Weyl group of Miller’s Lie algebra. Some of the types possess additional symmetries.
The power symmetry is a property of the ${}_{2}F_{1}$, ${}_{1}F_{1}$ and $F_{1}$ equation. For the $F_{1}$ equation it coincides with the basic symmetry.
The inversion symmetry is a feature of the ${}_{2}F_{1}$ equation, and connects the ${}_{1}F_{1}$ equation with the ${}_{2}F_{0}$ equation.
The ${}_{2}F_{1}$ and ${}_{2}F_{0}$ equations are invariant wrt swapping $a$ and $b$.
The Gegenbauer and Hermite equation are invariant wrt the symmetry $z\mapsto-z$.
The Lie algebra of the fourth column always contains the corresponding Miller’s Lie algebra (possibly, without the trivial term $\oplus{\mathbb{C}}$). By applying discrete symmetries from the fifth column to Miller’s Lie algebra, we can enlarge Miller’s Lie algebra to a subalgebra of the Lie algebra from the fourth column.
@@ 7 Hypergeometric polynomials
In the remaining subsections we return to the theory of hypergeometric class polynomials. We are convinced that the unified approach to their theory, which we present, based partly on [NU], has a considerable pedagogical value.
@@
The following well-known and easy proposition shows that hypergeometric class operators have many polynomial eigenfunctions. Actually, it is the property that characterizes this class among all second order differential operators.
@@
Let $\sigma(z)$, $\tau(z)$, $\mu(z)$ be arbitrary functions. Let $P_{n}(z)$, $n=0,1,2$, be polynomials such that $\deg P_{n}=n$ and $\eta_{n}\in{\mathbb{C}}$. Suppose that
$\left(\sigma(z)\partial_{z}^{2}+\tau(z)\partial_{z}+\mu(z)+\eta_{n}\right)P_{n}(z)=0,\ \ n=0,1,2.$
Then $\sigma(z)$ is a polynomial of degree at most $2$, $\tau(z)$ is a polynomial of degree at most 1 and $\mu(z)$ a polynomial of degree 00 (a complex number).
Note that for differential operators from Prop. 7.1, by replacing $\mu+\eta_{n}$ with $\eta_{n}$, without limiting the generality we can assume that $\mu=0$.
Let us describe some other simple facts about polynomial solutions of hypergeometric class equations:
@@
Let $\deg\sigma\leq 2$, $\deg\tau\leq 1$, $\eta\in{\mathbb{C}}$,
  1. 1.
    Suppose that $P$ is a polynomial of degree $n$ solving
    $\left(\sigma(z)\partial_{z}^{2}+\tau(z)\partial_{z}+\eta\right)P(z)=0.$ (7.1)
    Then
    $\begin{array}[]{l}n(n-1)\frac{\sigma^{\prime\prime}}{2}+n\tau^{\prime}+\eta=0.\end{array}$ (7.2)
  2. 2.
    If
    $k\frac{\sigma^{\prime\prime}}{2}+\tau^{\prime}\neq 0,\quad k\in{\mathbb{N}}_{0},$ (7.3)
    then the space of polynomial solutions of (7.1) is at most $1$-dimensional.
[…]
Suppose the space of polynomial solutions of (7.1) is 2-dimensional. We can assume that the degrees of these solutions are $0\leq n_{1}<n_{2}$. By (7.2),
@@
The following theorem shows how to construct hypergeometric class polynomials and describes some of their properties, which are easy to describe in a unified fashion. As usual, $\sigma$ denotes a polynomial of degree $\leq 2$, $\kappa$ a polynomial of degree $\leq 1$.
@@
    00 0 $=\Big{(}{{\cal H}}(\sigma,\kappa)+\omega\Big{)}P(z)$ (7.6)
    $=\Big{(}\sigma(z)\partial_{z}^{2}+(\sigma^{\prime}(z)+\kappa(z))\partial_{z}+\frac{\kappa^{\prime}}{2}+\omega\Big{)}P(z).$ (7.7)
@@
    00 0 $=\Big{(}{{\cal H}}(\sigma,\kappa)-n(n+1)\frac{\sigma^{\prime\prime}}{2}-(n+\frac{1}{2})\kappa^{\prime}\Big{)}P_{n}(\sigma,\kappa;z)$ (7.12)
@@
Proof.  1. and 2. are just a reformulations of Proposition 7.3, where instead of parameters $\tau,\eta$ we use $\kappa,\omega$.
To show 3. we note that the integral representation (7.11) satisfies the assumptions of Theorem 5.1 about Euler transforms. This representation implies both (7.12) and (7.10). $\Box$
(7.10) is usually called the Rodriguez formula.
Consider the descending ladder starting at $\kappa_{0}(z):=\kappa(z)$, $\omega_{0}=-\frac{\kappa^{\prime}}{2}$, that is,
@@
$\Big{(}{{\cal H}}(\sigma,\kappa_{-n})+\omega_{-n}\Big{)}P_{n}(\sigma,\kappa_{-n};z)$ $=0,$ (7.15)